OpenStudy (undeaddragon26):

Pizza Planet delivers pizza within a 15 mile radius of their store. What is the approximate area that they deliver to? Round to the nearest mile.

OpenStudy (undeaddragon26):

@Venny

OpenStudy (anonymous):

Area of a circle A=pi (r) ^2

OpenStudy (e.mccormick):

Or \(A=\pi r^2\) with the nifty math symbols.

OpenStudy (anonymous):

Wish I could do that :P

OpenStudy (undeaddragon26):

706.8583471?

OpenStudy (anonymous):

Yes that's what I got

OpenStudy (anonymous):

Don't forget to round :P

OpenStudy (anonymous):

round it and it's 707

OpenStudy (e.mccormick):

You can do it. Just type in \(\LaTeX\): `\(A=\pi r^2\)`

OpenStudy (undeaddragon26):

i did do it

OpenStudy (e.mccormick):

I meant Venny. Se said he wished he could type in \(\pi\), etc.

OpenStudy (anonymous):

Still don't get it, oh well